# Premerger phenomena in neutron-star binary coalescences

### v2

@@ VI.3 Resonant failure: luminosity and timescales
The resonant failure picture that has been alluded to throughout this Review is covered here and Sec. VI.4. The model was initially put forward by Tsang et al. 407 though has since been covered by many authors in different contexts. In a nutshell: a large amount of tidal energy can be rapidly siphoned into modes from the orbit when a resonance is hit depending on the respective overlap integral; see Sec. IV.1.3. For some modes, the resonant amplitude grows large enough to overstrain the crust (see Sec. III.1), releasing magnetoelastic energy that can theoretically fuel a gamma-ray flash. The way in this energy propagates is likely complicated depending on the local field strength, though the recent GR force-free electrodynamic simulations carried out by Most et al. 278 tracked Alfvén waves propagating out to the orbital light cylinder with crossing time $t_{\rm cross}^{2}\approx a^{3}/G(M_{A}+M_{B})$ where flares were formed self-consistently. For $a\approx 100$ km the light-crossing time is of order ms (see also Sec. VI.3.3). Although the mechanism responsible for the overstraining involves resonant pulsations rather than the gradual evolution of the magnetic field itself, the electromagnetic extraction of energy is totally analogous to that put forward for magnetar flares [399].
@@ VI.3.1 Energetics
Figure 16: Crustal strain pattern, $\sigma$ (equation 6), in the northern hemisphere as induced by a $g_{2}$-mode for an unmagnetised star with the SLy EOS in an equal mass binary with $M=1.27M_{\odot}$ and a constant stratification index $\delta=0.005$. For $\sigma_{\rm max}\lesssim 0.05$ (see Tab. 1) a complicated failure geometry will emerge. From Kuan et al. 231.
@@
More generally though, the maximum amount of energy during some window that can be liberated within a failed zone reads [e.g. 240],
@@
which is the collection of points such that the elastic maximum is exceeded at any given $t$.
@@
Under the approximations that (i) the energy released during a resonance timescale is just the integral of the energy density at the onset of resonance over the cracking area at the offset of resonance, and (ii) $u_{\text{rot}}$ is only associated with uniform rotation and frame dragging is ignorable (cf. Sec. IV.2.2), these energies were calculated by Kuan et al. 231 for some $g$-mode resonances (see Sec. VI.4.2). The main results are listed in Table 2 therein: for magnetar-level fields strengths, energy outputs can exceed $\sim 10^{46}$ erg each second. (How long such an event may last is covered in Sec. VI.3.2). This would be sufficient to power most precursors, noting in particular that only some small fraction of the crust actually fails (rather than global events) in those simulations; see Figure 16 for one such, $g_{2}$-strain pattern. If indeed the precursor was attributable to a $g_{2}$-mode, the isotropic luminosities quoted in Sec. V.4 would not be appropriate and in fact the energetics could easily be accommodated through expression (58).
@@ VI.3.2 Resonance window duration
which can be compared directly to observational durations from Table 3 assuming low redshift [cf. equation 1 in 446, and Sec. IV.2.2]. The (dimensionless) detuning parameter $\varepsilon$ can be calculated self-consistently with numerical simulations or the analytic approximation provided by equation (3.10) in Lai 233. It is generally expressible as
@@
for some parameter $1\lesssim\chi\lesssim\text{few}$. Although a value up to $\chi\approx 10$ was found for some $g$-modes by Kuan et al. 230, more realistic values in most cases are $\chi\sim\mathcal{O}(1)$. Either way, the numerical detuning parameter (64) can be somewhat larger than the analytic estimate made by Lai 233. More generally, $\chi$ depends on the EOS, mode family in question, spin, and microphysical (e.g. stratification) assumptions essentially because modes in stars with a specific mass can have the same frequency but different tidal overlap by tuning the aforementioned physics.
@@ VI.3.3 Launching timescale
One important aspect of the resonant-failure picture is that the flare will not be immediately launched once resonance, which is the quantity deducible from GWs for instance, is met. This couples in with the issues described in VI.1 for deducing precursor-merger waiting times. In particular, as detailed by Dichiara et al. 119, Kuan et al. 232, there are three timescales to consider. The first of these is the time it takes for overstraining to occur once resonance has been triggered. For the von Mises criterion (7), this failure happens instantaneously once expression (6) reaches a significantly large value. In a more realistic stress-strain model (e.g. Zhurkov; see Sec. III.2) there may be a short delay associated with this step; see Truesdell et al. 405.
The next timescale pertains to the delay for failure once an overstraining has been triggered from resonance. This was estimated by Tsang et al. 407 to be on the order of $\sim 1$ ms based on elastic-to-tidal energy ratios, though it is sensitive to both the overlap integral and mode frequency (see below equation 10 therein). These authors used a breaking-strain value of $\sigma_{\rm max}\sim 0.1$; all else being equal but using the Baiko and Chugunov 42 value $\sigma_{\rm max}\sim 0.04$ would increase this timescale by a factor $\sim(0.1/0.04)^{2}\approx 6$ (see Tab. 1). Higher eigenfrequencies or energies would reduce it.
The final timescale of relevance, $t_{\rm emit}$, relates to how long it takes for emissions to be generated following the failure [see 398, 399, regarding the magnetar flare context]. Neill et al. 287 argue $t_{\rm emit}$ can be as long as $\sim 0.1$ s for $B\sim 10^{13}$ G — see equation (12) therein — again assuming $\sigma_{\rm max}\sim 0.1$ from Horowitz and Kadau 193. The result scales like [232]
@@
For most precursors, $t_{\rm emit}$ will be of order tens of ms. In support of this, Most et al. 278 suggest $t_{\rm emit}$ is related to a light crossing-time that is relatively short unless $a\gg 100$ km.
Altogether there are thus eight timescales to keep track of when trying to directly compare a tidal resonance to a precursor timing observation, at least in the resonance picture. These are the five detailed in Sec. VI.1 relating to jet generation and breakout, and the three described above (or only one in the case of magnetospheric interactions). This highlights the non-trivial nature of the problem, where future numerical simulations of jet break out and radiation transport are likely to prove particularly powerful.
@@ VI.4 Resonant failure: some important families
Figure 17: Resonant amplitudes (expressed via energies in units of $E_{0}=GM_{A}^{2}/R_{A}$) obtained for the crust-core interface modes $i_{2}$ for a polytropic star in Newtonian gravity ($\Gamma_{1}$ is our $\gamma$), where elastic terms in the crust are self-consistently included for a variety of shear modulii $[\tilde{\mu}$; normalised according to equation 73 in Passamonti et al. 302]. Circles (squares) denote distances associated with the moment of crustal failure, if applicable, for a stronger (weaker) breaking strain of 0.1 (0.04) under a von Mises criterion. From Passamonti et al. 302 with permission.
@@
As a neutron star cools, it is expected to undergo a number of localised phase transitions. These can relate to superfluidity and conductivity (Sec. III.4), and the formation of a crust and ocean layers (Sec. III.1). The presence of these solid-liquid discontinuities allows for a family of interface ($i$-) modes to exist. These were the first family that were considered viable for breaking and arguably remain the strongest candidate [407, 406, 119], at least for the crust-core ($i_{2})$ rather than crust-ocean ($i_{1}$) variety, though the latter can also induce non-trivial strain and crustal failure (see Sec. VI.4.3). This is because the overlap integrals tend to be large and $i$-mode eigenfrequencies are extremely sensitive to the transition density and stellar macrophysics, allowing for many events to be explained by this one mode in different stars [287]. Figure 17 shows predictions for the (appropriately normalised) energies deposited into $i$-modes, as found by Passamonti et al. 302, who self-consistently included elastic terms into their (Newtonian) equations of motion.
@@
Passamonti et al. 302 found $i$-mode overlaps to be two orders of magnitude weaker than that found by Tsang et al. 407. Although still plausibly large enough to break the crust (at least for sufficiently large shear modulii; see Fig. 17), this demonstrates that $i$-modes may not be generically optimal for any given neutron star configuration when incorporating elastic aspects. Another possible origin for this large discrepancy is rooted in the mixed use of Newtonian overlaps/perturbations but with a GR star. Although the use of such a hybrid scheme to solve for stellar spectrum is common in the literature [and was used by 407, for instance], it leads to some non-orthogonality between modes and there can then be “leakage” of one mode to others when computing overlap integrals [^8] In GR, modes are not strictly orthogonal but the inner product between them is much smaller than the extent of this leakage; see also Footnote 6. [see Appendix A of 273, for a qualitative discussion].
Overall, however, owing to the sensitivity of $i$-modes to crustal microphysics, these modes represent a powerful probe of stellar structure [e.g. for the possible existence of pasta phases; 287, 360].
@@
As described throughout, these modes are due to buoyancy resulting from thermal or composition gradients [including discontinuities; 141, 261, 361] sourced intrinsically or through the accretion of matter from a companion, crustal movements, or magnetic interactions [e.g. 125]. For cold stars in a late-stage binary, the frequencies of (normal-fluid) $g$-modes can range from tens to hundreds of Hz, indicating resonances occur $1-10$ s prior to merger when spin is insignificant (see Sec. IV.3). Interestingly, maximum amplitudes and frequency tend to increase with stellar mass for the $g_{1}$-modes, while the amplitudes tend to decrease for $g_{2}$-modes; see Table 1 in Suvorov et al. 386. On the other hand, $g$-modes in less compact stars have longer growth times, compensating their weaker coupling to the tidal field to some extent. As such, the maximally induced stress exhibits a parabola-like shape, at least when the stratification is held fixed, as seen in Figure 18. The local minima of these curves has been referred to as “tidal neutrality”, as described in Appendix A of Kuan et al. 228.
@@
Figure 18: Dependence of the maximally induced stress (equation 6) due to $g_{1}$- modes for various EOS (same for both stars with $q=1$) as functions of the stellar compactness $\mathcal{C}=M_{1.4}/R_{10\text{km}}$. The grey dashed lines represents the von Mises criterion identified by Baiko and Chugunov 42 (lower) and Horowitz and Kadau 193 (upper). A fixed value of $\delta=0.005$ is used. From Kuan et al. 231.
@@
In addition to the intrinsic mode properties, the binary mass-ratio also influences the resonance time-scale through the tidal potential (13), and hence the maximal stress also. Fig. 19 demonstrates that a larger strain is generally manifest when the companion is heavier (i.e., the $g_{1}$-mode issues a greater $\sigma$) for a given chirp mass, $\mathcal{M}=M_{A}q^{3/5}/(1+q)^{1/5}$, of the binary, which can be precisely estimated with GW analysis [e.g., 102]. This occurs because when $\mathcal{M}$ is fixed a larger $q$ implies a lower mass and pushes the system into the left-half of the compactness range considered in Figure 18, noting the use of WFF1 (Sec. II.1.1). For other EOS, the trend may reverse.
@@
Figure 19: Sequences of maximal strain (redder shades indicating a greater value for $\sigma$) driven by $g_{1}$-mode resonances for the WFF1 EOS for various chirp masses $\mathcal{M}$ and mass-ratios. The blue vertical line shows the chirp mass of the progenitor of GW 170817, while the red corresponds to GW 190425. From Kuan et al. 231.
@@
As discussed in IV.3, spin reduces the frequency of retrograde modes. When the stellar angular momentum is in the opposite direction to that of the orbit, the oscillation rate of tidally-susceptible modes (whose phase pattern velocity matches the orbit) is slowed, pushing the resonance time earlier. Although perhaps counter-intuitive, Lai 233 showed that the scaling of the saturation amplitude after resonance is $\propto\omega_{\alpha}^{-5/6}$ under the stationary phase approximation (i.e., when the orbital decay rate is much slower than the mode frequency): an earlier resonance thus generally renders a higher amplitude, leading to the general increase with maximal strain as a function of spin as presented in Figure 20 for fixed overlap. In principle, this increase reaches a peak at the spin such that the mode will be likely smeared out by chemical reaction, leading to an abrupt zeroing of the strain. Note, however, such a smearing was not modelled by Kuan et al. 231: the reason for the drop in $\sigma$ in this case is because the mode frequency drops enough that resonance window falls outside of the computational domain at some minimum $\Omega_{\rm Orb}(t=0)$.
@@
Although $g$-modes reside mostly in the core, they are capable of exerting strain at the bottom of crust for two reasons. (i) They penetrate into the crust when $\delta$ is sufficiently large; in fact, Passamonti et al. 302, Gittins and Andersson 158 demonstrated that the $g_{1}$-mode eigenfunction can resemble that of the $i$-mode though the frequencies differ [cf. 263]. (ii) Their tangential motion is not necessarily small even when buoyancy is somewhat quenched by the shear modulus. Should some crustal activities be triggered by $g$-modes and lead to observables, information about the internal stratification, thermal properties, cooling, phase transitions [including gravitational ones; 229], and heat transport in neutron stars could all be gleaned.
@@
Figure 20: Similar to Fig. 18 but as functions of the stellar spin for fixed compactness. Plot legends for the systems define the mutually-abiding EOS together with individual masses in units of $M_{\odot}$. Sharp drops after a certain spin for different modes results because the mode frequency drops lower than the initial tidal pushing frequency used in the numerical computation, and may be analogous to the spectral “washing out” described in Sec. III.3 once the mode frequencies drop too low. From Kuan et al. 231.

### v3

@@ VI.3 Resonant failure: luminosity and timescales
The resonant failure picture that has been alluded to throughout this Review is covered here and Sec. VI.4. The model was initially put forward by [421] though has since been covered by many authors in different contexts. In a nutshell: a large amount of tidal energy can be rapidly siphoned into modes from the orbit when a resonance is hit depending on the respective overlap integral; see Sec. IV.1.3. For some modes, the resonant amplitude grows large enough to overstrain the crust (see Sec. III.1), releasing magnetoelastic energy that can theoretically fuel a gamma-ray flash. The way in this energy propagates is likely complicated depending on the local field strength, though the recent GR force-free electrodynamic simulations carried out by [290] tracked Alfvén waves propagating out to the orbital light cylinder with crossing time $t_{\rm cross}^{2}\approx a^{3}/G(M_{A}+M_{B})$ where flares were formed self-consistently. For $a\approx 100$ km the light-crossing time is of order ms (see also Sec. VI.3.3). Although the mechanism responsible for the overstraining involves resonant pulsations rather than the gradual evolution of the magnetic field itself, the electromagnetic extraction of energy is totally analogous to that put forward for magnetar flares [413].
@@ VI.3.1 Energetics
Figure 16: Crustal strain pattern, $\sigma$ (equation 6), in the northern hemisphere as induced by a $g_{2}$-mode for an unmagnetised star with the SLy EOS in an equal mass binary with $M=1.27M_{\odot}$ and a constant stratification index $\delta=0.005$. For $\sigma_{\rm max}\lesssim 0.05$ (see Tab. 1) a complicated failure geometry will emerge. From Kuan et al. 240.
@@
More generally though, the maximum amount of energy during some window that can be liberated within a failed zone reads [e.g. 249],
@@
which is just the set where the elastic maximum is exceeded at any given $t$.
@@
Under the approximations that (i) the energy released during a resonance timescale is just the integral of the energy density at the onset of resonance over the cracking area at the offset of resonance, and (ii) $u_{\text{rot}}$ is only associated with uniform rotation and frame dragging is ignorable (cf. Sec. IV.2.2), these energies were calculated by Kuan et al. 240 for some $g$-mode resonances (see Sec. VI.4.2). The main results are listed in Table 2 therein: for magnetar-level fields strengths, energy outputs can exceed $\sim 10^{46}$ erg each second. (How long such an event may last is covered in Sec. VI.3.2). This would be sufficient to power most precursors, noting in particular that only some small fraction of the crust actually fails (rather than global events) in those simulations; see Figure 16 for one such, $g_{2}$-strain pattern. If indeed the precursor was attributable to a $g_{2}$-mode, the isotropic luminosities quoted in Sec. V.4 would not be appropriate and in fact the energetics could easily be accommodated through expression (58).
@@ VI.3.2 Resonance window duration
which can be compared directly to observational durations from Table 3 assuming low redshift [cf. equation 1 in 461, and Sec. IV.2.2]. The (dimensionless) detuning parameter $\varepsilon$ can be calculated self-consistently with numerical simulations or the analytic approximation provided by equation (3.10) in Lai 242. It is generally expressible as
@@
for some parameter $1\lesssim\chi\lesssim\text{few}$. Although a value up to $\chi\approx 10$ was found for some $g$-modes by Kuan et al. 239, more realistic values in most cases are $\chi\sim\mathcal{O}(1)$. Either way, the numerical detuning parameter (64) can be somewhat larger than the analytic estimate made by Lai 242. More generally, $\chi$ depends on the EOS, mode family in question, spin, and microphysical (e.g. stratification) assumptions essentially because modes in stars with a specific mass can have the same frequency but different tidal overlap by tuning the aforementioned physics.
@@ VI.3.3 Launching timescale
One important aspect of the resonant-failure picture is that the flare will not be immediately launched once resonance, which is the quantity deducible from GWs for instance, is met. This couples in with the issues described in VI.1 for deducing precursor-merger waiting times. In particular, as detailed by Dichiara et al. 119, Kuan et al. 241, there are three timescales to consider. The first of these is the time it takes for overstraining to occur once resonance has been triggered. For the von Mises criterion (7), this failure happens instantaneously once expression (6) reaches a significantly large value. In a more realistic stress-strain model (e.g. Zhurkov; see Sec. III.2) there may be a short delay associated with this step; see Truesdell et al. 419.
The next timescale pertains to the delay for failure once an overstraining has been triggered from resonance. This was estimated by Tsang et al. 421 to be on the order of $\sim 1$ ms based on elastic-to-tidal energy ratios, though it is sensitive to both the overlap integral and mode frequency (see below equation 10 therein). These authors used a breaking-strain value of $\sigma_{\rm max}\sim 0.1$; all else being equal but using the Baiko and Chugunov 42 value $\sigma_{\rm max}\sim 0.04$ would increase this timescale by a factor $\sim(0.1/0.04)^{2}\approx 6$ (see Tab. 1). Higher eigenfrequencies or energies would reduce it.
The final timescale of relevance, $t_{\rm emit}$, relates to how long it takes for emissions to be generated following the failure [see 412, 413, regarding the magnetar flare context]. Neill et al. 300 argue $t_{\rm emit}$ can be as long as $\sim 0.1$ s for $B\sim 10^{13}$ G — see equation (12) therein — again assuming $\sigma_{\rm max}\sim 0.1$ from Horowitz and Kadau 200. The result scales like [241]
@@
For most precursors, $t_{\rm emit}$ will be of order tens of ms. In support of this, Most et al. 290 suggest $t_{\rm emit}$ is related to a light crossing-time that is relatively short unless $a\gg 100$ km.
Altogether there are thus eight timescales to keep track of when trying to directly compare a tidal resonance to a precursor timing observation, at least in the resonance picture for bright nonthermal precursors. These are the five detailed in Sec. VI.1 relating to jet generation and breakout, and the three described above (or only one in the case of magnetospheric interactions). This highlights the non-trivial nature of the problem, where future numerical simulations of jet break out and radiation transport are likely to prove particularly powerful. In fact, for thermal or comparatively low luminosity precursors, an additional set of physical timescales become relevant relating to collisions of fireball shells. As described by Neill et al. [300], regions surrounding the emission site may be optically thick to pair production and multiple pair-photon fireballs could be launched which, upon collision, create shocks and synchrotron emissions together with inherent blackbody emissions with luminosities reaching $\sim 10^{48}$ erg/s. The collision times of these shells depends on the Lorentz factors set through the degree of mass loading, with theoretical ranges spanning a few orders of magnitude suggesting a wide range of viable timescales.
@@ VI.4 Resonant failure: some important families
Figure 17: Resonant amplitudes (expressed via energies in units of $E_{0}=GM_{A}^{2}/R_{A}$) obtained for the crust-core interface modes $i_{2}$ for a polytropic star in Newtonian gravity ($\Gamma_{1}$ is our $\gamma$), where elastic terms in the crust are self-consistently included for a variety of shear modulii $[\tilde{\mu}$; normalised according to equation 73 in Passamonti et al. 315]. Circles (squares) denote distances associated with the moment of crustal failure, if applicable, for a stronger (weaker) breaking strain of 0.1 (0.04) under a von Mises criterion. From Passamonti et al. 315 with permission.
@@
As a neutron star cools, it is expected to undergo a number of localised phase transitions. These can relate to superfluidity and conductivity (Sec. III.4), and the formation of a crust and ocean layers (Sec. III.1). The presence of these solid-liquid discontinuities allows for a family of interface ($i$-) modes to exist. These were the first family that were considered viable for breaking and arguably remain the strongest candidate [421, 420, 119], at least for the crust-core ($i_{2})$ rather than crust-ocean ($i_{1}$) variety, though the latter can also induce non-trivial strain and crustal failure (see Sec. VI.4.3). This is because the overlap integrals tend to be large and $i$-mode eigenfrequencies are extremely sensitive to the transition density and stellar macrophysics, allowing for many events to be explained by this one mode in different stars [300]. Figure 17 shows predictions for the (appropriately normalised) energies deposited into $i$-modes, as found by [315], who self-consistently included elastic terms into their (Newtonian) equations of motion.
@@
Passamonti et al. [315] found $i$-mode overlaps to be two orders of magnitude weaker than that found by Tsang et al. [421]. Although still plausibly large enough to break the crust (at least for sufficiently large shear modulii; see Fig. 17), this demonstrates that $i$-modes may not be generically optimal for any given neutron star configuration when incorporating elastic aspects. Another possible origin for this large discrepancy is rooted in the mixed use of Newtonian overlaps/perturbations but with a GR star. Although the use of such a hybrid scheme to solve for stellar spectrum is common in the literature (and was used by [421] for instance), it leads to some non-orthogonality between modes and there can then be “leakage” of one mode to others when computing overlap integrals [^8] In GR, modes are not strictly orthogonal but the inner product between them is much smaller than the extent of this leakage; see also Footnote 6. [see Appendix A of [284] for a qualitative discussion].
Overall, however, owing to the sensitivity of $i$-modes to crustal microphysics, these modes represent a powerful probe of stellar structure [e.g. for the possible existence of pasta phases; 300, 374]. Neill et al. [299] demonstrated that constraints for the nuclear symmetry parameters — namely the (isovector) symmetry energy parameters $J$, $L$, and $K_{\rm sym}$ — of a (nucleonic) EOS can be made when combining precursor measurements, assuming $i$-mode triggers, with GWs. The main results are described in figure 3 therein, illustrating the complementary nature of such constraints with those obtained from terrestrial (e.g. PREX-II [345], where the neutron skin thickness of the heaviest lead isotopes were measured) and astrophysical experiments. For example, for a given $i$-mode frequency and ignoring the impacts of rotation or magnetic fields, shifts of $\sim 20$ MeV in uncertainty ranges for $L$ could be anticipated.
There are open issues in the modelling of $i$-modes however, such as relates to consistent GR treatments, the choice of shear modulus profile, and the possibility of anisotropic nuclear clusters and pasta. Accurately incorporating these elements is a difficult task owing to both conceptual and numerical issues, and it is unclear how much anticipated constraints on parameter spaces may vary.
@@
As described throughout, these modes are due to buoyancy resulting from thermal or composition gradients [including discontinuities; 144, 271, 375] sourced intrinsically or through the accretion of matter from a companion, crustal movements, or magnetic interactions [e.g. 126]. For cold stars in a late-stage binary, the frequencies of (normal-fluid) $g$-modes can range from tens to hundreds of Hz, indicating resonances occur $1-10$ s prior to merger when spin is insignificant (see Sec. IV.3). Interestingly, maximum amplitudes and frequency tend to increase with stellar mass for the $g_{1}$-modes, while the amplitudes tend to decrease for $g_{2}$-modes; see Table 1 in Suvorov et al. 400. On the other hand, $g$-modes in less compact stars have longer growth times, compensating their weaker coupling to the tidal field to some extent. As such, the maximally induced stress exhibits a parabola-like shape, at least when the stratification is held fixed, as seen in Figure 18. The local minima of these curves has been referred to as “tidal neutrality”, as described in Appendix A of Kuan et al. 237.
@@
Figure 18: Dependence of the maximally induced stress (equation 6) due to $g_{1}$- modes for various EOS (same for both stars with $q=1$) as functions of the stellar compactness $\mathcal{C}=M_{1.4}/R_{10\text{km}}$. The grey dashed lines represents the von Mises criterion identified by Baiko and Chugunov 42 (lower) and Horowitz and Kadau 200 (upper). A fixed value of $\delta=0.005$ is used. From Kuan et al. 240.
@@
In addition to the intrinsic mode properties, the binary mass-ratio also influences the resonance time-scale through the tidal potential (13), and hence the maximal stress also. Fig. 19 demonstrates that a larger strain is generally manifest when the companion is heavier (i.e., the $g_{1}$-mode issues a greater $\sigma$) for a given chirp mass, $\mathcal{M}=M_{A}q^{3/5}/(1+q)^{1/5}$, of the binary, which can be precisely estimated with GW analysis [e.g., 101]. This occurs because when $\mathcal{M}$ is fixed a larger $q$ implies a lower mass and pushes the system into the left-half of the compactness range considered in Figure 18, noting the use of WFF1 (Sec. II.1.1). For other EOS, the trend may reverse.
@@
Figure 19: Sequences of maximal strain (redder shades indicating a greater value for $\sigma$) driven by $g_{1}$-mode resonances for the WFF1 EOS for various chirp masses $\mathcal{M}$ and mass-ratios. The blue vertical line shows the chirp mass of the progenitor of GW 170817, while the red corresponds to GW 190425. From Kuan et al. 240.
@@
As discussed in Sec. IV.3, spin reduces the frequency of retrograde modes. When the stellar angular momentum is in the opposite direction to that of the orbit, the oscillation rate of tidally-susceptible modes (whose phase pattern velocity matches the orbit) is slowed, pushing the resonance time earlier. Although perhaps counter-intuitive, Lai 242 showed that the scaling of the saturation amplitude after resonance is $\propto\omega_{\alpha}^{-5/6}$ under the stationary phase approximation (i.e., when the orbital decay rate is much slower than the mode frequency): an earlier resonance thus generally renders a higher amplitude, leading to the general increase with maximal strain as a function of spin as presented in Figure 20 for fixed overlap. In principle, this increase reaches a peak at the spin such that the mode will be likely smeared out by chemical reaction, leading to an abrupt zeroing of the strain. Note, however, such a smearing was not modelled by [240]: the reason for the drop in $\sigma$ in this case is because the mode frequency drops enough that resonance window falls outside of the computational domain at some minimum $\Omega_{\rm Orb}(t=0)$.
@@
Although $g$-modes reside mostly in the core, they are capable of exerting strain at the bottom of crust for two reasons. (i) They penetrate into the crust when $\delta$ is sufficiently large; in fact, Passamonti et al. 315, Gittins and Andersson 163 demonstrated that the $g_{1}$-mode eigenfunction can resemble that of the $i$-mode though the frequencies differ [cf. 273]. (ii) Their tangential motion is not necessarily small even when buoyancy is somewhat quenched by the shear modulus. Should some crustal activities be triggered by $g$-modes and lead to observables, information about the internal stratification, thermal properties, cooling, phase transitions [including gravitational ones; 238], and heat transport in neutron stars could all be gleaned.
There are open problems in this direction related to the handling of the microphysics that more sophisticated models hope to treat in future. Even with respect to models described above that do include crustal elasticity explicitly, compositional stratifications at the base of the crust may be driven by neutron drip from nuclear clusters, which is likely to occur at timescales faster than that of g-mode oscillations, thereby modifying the mode eigenfunctions there relative to models where perturbed adiabatic indices were put in by hand. Other models described throughout that do use tabulated EOS data for describing the crust and boundary layer composition (see Fig. 5) but which do not include crustal elasticity may similarly overestimate the crustal strain. Efforts to include all such components for $g$-modes are in progress.
@@
Figure 20: Similar to Fig. 18 but as functions of the stellar spin for fixed compactness. Plot legends for the systems define the EOS that is obeyed by the primary and the companion together with individual masses in units of $M_{\odot}$. Note the sharp drops after a certain spin for different modes results because the mode frequency drops lower than the initial tidal pushing frequency used in the numerical computation, and may be analogous to the spectral “washing out” described in Sec. III.3 once the mode frequencies drop too low. From Kuan et al. 240.